A survey on subjects being taken by 250 college students in La Union revealed the following information: (Subject-number of students taking the subject) Mathematics - 145; Filipino - 90; English - 88; Math and Filipino - 25; Filipino and English - 38; Math and English - 59; Math, Filipino and English - 15. How many did not take any of the three subjects?
1
Expert's answer
2021-10-15T10:34:33-0400
"N(U)=250, N(M)=145, N(F)=90, N(E)=88,"

"N(M\cap F)=25, N(M\cap E)=59, N(F\cap E)=38,"

"N(M\cap F\cap E)=15"

"N(M\cup F\cup E)=N(M)+N(F)+N(E)"

"-N(M\cap F)-N(M\cap E)-N(F\cap E)"

"+N(M\cap F\cap E)=145+90+88"

"-25-59-38+15=216"

"N(\overline{M\cup F\cup E})=N(U)-N(M\cup F\cup E)"

"=250-216=34"

34 students did not take any of the three subjects.